OpenStudy (anonymous):

integration of (1+e^x)^2 /e^x

OpenStudy (amistre64):

expand the top, divide out the denominator and see what we got left?

OpenStudy (anonymous):

\[\int\limits_{a}^{b} ((1+e^x)^2)/e^x \] hmmm... long time since Cal II. Maybe Integration by Parts?

OpenStudy (anonymous):

\[\int \left(e^{-x} + e^{x} + 1 \right)dx = -e^{-x} + e^x + x \]

OpenStudy (amistre64):

\[\int\frac{(1+e^x)^2}{e^x}dx\] \[\int\frac{1+2e^x+e^{2x}}{e^x}dx\] \[\int\frac{1}{e^x}+\frac{2e^x}{e^x}+\frac{e^{2x}}{e^x}dx\] \[\int e^{-x}+{2e^x}+{e^x}dx\]

OpenStudy (amistre64):

typoed myself lol

OpenStudy (amistre64):

\[\int e^{-x}+2+{e^x}dx\] as i see it

OpenStudy (anonymous):

yup it should be 2 sorry my bad ..amistre is right

OpenStudy (anonymous):

Whoops I made a mistake.

OpenStudy (amistre64):

my brain did it right, and my fingers argued with me lol

OpenStudy (anonymous):

so the final answer is \[-e^-x+2x+e^x ?\]